# A short program in MuPAD that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$

### v30

@@ Abstract
1 Introduction
@@
(cf. [9, pp. 233–235]). A computation in the limit of a function $f:\mathbb{N}\to\mathbb{N}$ is a semi-algorithm which takes as input a non-negative integer $n$ and for every $m\in\mathbb{N}$ prints a non-negative integer $\xi(n,m)$ such that $\lim\limits_{m\to\infty}\xi(n,m)=f(n)$.
It is known that there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which is not computable, see Theorem 1. Every known proof of this fact does not lead to the existence of a short computer program that computes $f$ in the limit. In particular, this observation applies to the proof of Theorem 1 in [8], see also Observation 1.
Observation 1.
Let $\varphi$ be a computable bijection from $\mathbb{N}$ to the set of all Diophantine equations. For $n\in\mathbb{N}$, let
$\theta(n)=\left\{\begin{array}[]{cl}1,&\text{if~}\varphi(n)\text{~is~solvable~in~non-negative~integers}\\ 0,&\text{otherwise}\end{array}\right$.
The function $\theta:\mathbb{N}\to\mathbb{N}$ is computable in the limit. A negative solution to Hilbert’s 10th problem implies that the function $\theta$ is not computable. There is no known $\varphi$ for which there exists a short computer program that computes $\theta$ in the limit.
MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b. In this article, we present a short program in MuPAD that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
@@
([8, p. 118]). There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
@@ 3 A short program in MuPAD that computes $f$ in the limit
It follows from Lemmas 1 and 2. ∎
@@
Observation 4.
For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in\{0,\ldots,m\}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
@@
([1], [2], [3], [4, pp. 341–342], [6, p. 42], [7, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
@@ 5 A short program in MuPAD that computes $\beta$ in the limit
It follows from Lemmas 3 and 4. ∎

### v31

# A short program that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$

@@ Abstract
1 The Collatz problem leads to a short program that computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$ of unknown computability
@@
(cf. [10, pp. 233–235]). A computation in the limit of a function $f:\mathbb{N}\to\mathbb{N}$ is a semi-algorithm which takes as input a non-negative integer $n$ and for every $m\in\mathbb{N}$ prints a non-negative integer $\xi(n,m)$ such that $\lim\limits_{m\to\infty}\xi(n,m)=f(n)$.
By Definition 1, a function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit when there exists an infinite computation which takes as input a non-negative integer $n$ and prints a non-negative integer on each iteration and prints $f(n)$ on each sufficiently high iteration.
It is known that there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which is not computable, see Theorem 1. Every known proof of this fact does not lead to the existence of a short computer program that computes $f$ in the limit. So far, short computer programs can only compute in the limit functions from $\mathbb{N}$ to $\mathbb{N}$ whose computability is proven or unknown.
Lemma 1.
For every $n\in\mathbb{N}$,
$\frac{\textstyle\text{sign}(n-1)\cdot(2n+(1-(-1)^{n})\cdot(5n+2))}{\textstyle 4}=\left\{\begin{array}[]{rl}0{\rm,}&{\rm if~}n=1\\ \frac{\textstyle n}{\textstyle 2}{\rm,}&{\rm if~}n{\rm~is~even}\\ 3n+1{\rm,}&{\rm if~}n{\rm~is~odd~and~}n\neq 1\end{array}\right$.
MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b. By Lemma 1, the following program in MuPAD computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$.
input("Input a non-negative integer n",n):
while TRUE do
print(sign(n)):
n:=sign(n-1)*(2*n+(1-(-1)^n)*(5*n+2))/4:
end_while:
The computability of $\gamma$ is unknown, see [1, p. 79]. The Collatz conjecture implies that $\gamma(n)=0$ for every $n\in\mathbb{N}$.
@@
([9, p. 118]). There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
@@ 3 A short program in MuPAD that computes $f$ in the limit
It follows from Lemmas 2 and 3. ∎
@@
For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the smallest $b\in\{0,\ldots,m\}$ such that every tuple $(x_{0},\ldots,x_{n})\in\{0,\ldots,m\}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
@@
([2], [3], [4], [5, pp. 341–342], [7, p. 42], [8, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
@@ 5 A short program in MuPAD that computes $\beta$ in the limit
It follows from Lemmas 4 and 5. ∎